# OptiCarVis: Improving Automated Vehicle Functionality Visualizations Using Bayesian Optimization to Enhance User Experience

### v1

@@ Abstract.
journalyear: 2025copyright: rightsretaineddoi: XXXXXXX.XXXXXXXsubmissionid: 4661ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing Empirical studies in visualizationccs: Human-centered computing Empirical studies in HCI
@@
The figure shows the OptiCarVis flow chart. The Bayesian Optimization process is applied to the results, which are repeatedly evaluated by study participants via a questionnaire.
@@ 1. Introduction
However, user acceptance of AV technology is not guaranteed. Studies have shown that many potential users are concerned about AV reliability (Schoettle and Sivak, 2014; Kyriakidis et al., 2015), which refers to AVs’ ability to perform driving tasks safely across different situations consistently. Both undertrust (e.g., leading to not using AVs) and overtrust (e.g., inadequately supervising the operation) present challenges for AV usage if the user’s trust is inappropriate to the actual AV reliability (Lee and See, 2004). Therefore, previous works have shown that visualizations of AV functionality are a way to enhance user experience (Colley et al., 2022b; Flohr et al., 2023; Schneider et al., 2021). Suggested visualization solutions to overcome undertrust highlighted other road users in foggy scenarios (Winter et al., 2019). In the Connected Automated Driving (CAD) context, proposals include visualizing external sensor coverage (henceforth called “CAD-covered area”) and detected road users occluded, for example, by a building (henceforth called “occluded cars”) (Müller et al., 2022). Regarding overtrust, visualizing the internal functionalities of AVs (e.g., Situation Detection, Situation Prediction, or Trajectory Planning) and their inherent uncertainty were evaluated (Colley et al., 2020, 2021a, 2022b; Flohr et al., 2023).
@@
However, the diversity of passengers complicates the design of AV functionality visualizations. Passengers’ subjective perceptions of safety, trust, and aesthetics differ (Mara and Meyer, 2022). In addition, passengers’ understanding of AVs’ internal functionality depends on individual knowledge and attitudes towards technology (Zhang et al., 2023). Thus, to support user understanding, designers are tasked with balancing various design objectives within a complex design space (i.e., the set of possible design parameter values), for example, determining the size, transparency, and necessity of visualization elements. Traditional design methods relied on the user-centered design process (see ISO 9241 (International Organization for Standardization, 2019)), standards (e.g., ISO 15005 (International Organization for Standardization, 2017)), and guidelines (e.g., the JAMA Guidelines for In-vehicle Display System (Yukinobu Nakamura, 2008)) but also experience, trial and error, and intuition especially when designing for novel experiences as with AVs, resulting in resource-intensive user evaluations.
To align visualization designs with passengers’ diverse needs and preferences, previous work explored personalization approaches. For instance, Normark (2015) enabled passengers to manually personalize icons’ size, location, and color on the dashboard, center stack, and Head-Up Display (HUD). However, without design experience, it can be challenging to translate personal needs and preferences into design parameter values. In contrast, computational methods have been used to help designers choose a visualization design. For example, Zhong et al. (2023) elicited weights to designers’ ratings, which informed a decision model to choose a design out of three candidate HUD design schemes. Likewise, Yunuo et al. (2022) obtained weighted ratings by drivers and, accordingly, their model selected a HUD design out of 18 candidates. However, these previous works did not explore the design space with continuous values but only used a small set of pre-defined visualizations (see (Yunuo et al., 2022; Zhong et al., 2023)), which could potentially miss designs that better meet passengers’ needs and preferences. Besides, while these works measure the usability (Yunuo et al., 2022; Zhong et al., 2023), they disregard perceived safety and trust, which are crucial for user acceptance of AVs (Adnan et al., 2018) and, ultimately, the user experience.
@@
If a design should enhance user experience, involving humans in the design process is important to ensure alignment with their needs and preferences. Therefore, a Human-in-the-Loop (HITL) process (see (Chan et al., 2022)) could be used that iteratively presents users with design variants. Their feedback is used in each iteration to optimize the design parameter values to meet the users’ needs and preferences better. However, such an optimization is challenging because of multiple design objectives (e.g., increasing perceived safety and trust). In non-automotive User Interface (UI) domains, approaches using (Multi-Objective) Bayesian Optimization (BO and MOBO) have been emerging to solve design optimization problems (Chan et al., 2022; Liao et al., 2023; Chandramouli et al., 2023; Koyama and Goto, 2022; Kadner et al., 2021). BO optimizes designs by predicting which changes in the parameter values will most effectively meet the design objective. In MOBO–an extension of BO that can handle multiple objectives–trade-offs among potentially conflicting objectives are managed by identifying points on the Pareto front, where any improvement in one objective would lead to another objective’s deterioration. For instance, on this front, enhancing perceived safety by showing more visualizations may increase cognitive load. Hereafter, one such design is referred to as Pareto-optimal (Marler and Arora, 2004).
@@
We demonstrate OptiCarVis by designing visualizations of an AV’s functional levels (Situation Detection, Situation Prediction, and Trajectory Planning), as well as the CAD-covered area and occluded cars, and the general information (i.e., AV speed, AV destination, and current time) on a HUD, as these can convey relevant information on various stages of the automated driving task (Colley et al., 2022b; Müller et al., 2022; Flohr et al., 2023). The visualizations include semantic segmentation for Situation Detection, pedestrian intention icons and other road users’ trajectories for Situation Prediction, and a trajectory for the AV’s own Trajectory Planning. We created an AV driving environment in Unity, where end-users view these visualizations on a simulated Augmented Reality (AR) Windshield Display (WSD) shown on their computer screen. In each HITL iteration (see Figure 1), the MOBO first estimates parameter values in the design space of functionality visualizations. End-users then experience the chosen values and provide subjective ratings after the ride. Next, the MOBO uses these ratings to understand how well the design has met the objectives and optimizes the parameter values for the subsequent iteration. This is repeated until design objectives are met (e.g., maximum trust on the rating scale).
@@
To quantify the optimization’s effectiveness, end-users evaluate their final design after the condition (regarding the cognitive load, predictability, trust, perceived safety, usefulness, satisfaction, and aesthetics). We also compare the HITL MOBO results to the baseline crafted by automotive UI experts (N=8). Besides, we capture end-user behaviors through webcam-based eye-tracking to monitor their reactions to MOBO-driven visualization designs and verify their attention throughout the study.
We conducted a between-subject online study with 117 participants. The study found that OptiCarVis returned personalized design parameters for visualizing AVs’ Situation Detection, Situation Prediction, and own Trajectory Planning that significantly enhanced perceived safety relative to participants’ custom design, improved predictability compared to the pre-defined expert design, and increased trust, usefulness, and satisfaction over both. Participants’ feedback expressed satisfaction with the visualization design process, indicating a sense of involvement. Many deemed the design optimal but suggested incorporating additional visualization elements and expanding driving scenarios.
By leveraging OptiCarVis, automotive UI designers and end-users can navigate complex design spaces, potentially resulting in more passenger-centric UIs that could significantly increase their perceived safety, trust, and acceptance of AVs. Moreover, the direct integration of end-users’ feedback into the design could inspire the development of more cooperative and effective end-user optimization methods that are implicitly integrated into future automotive UIs.
Contribution Statement: (1) OptiCarVis—the computational optimization of AV functionality visualization design using HITL MOBO to improve end-user trust, perceived safety, acceptance, and understanding of AVs and decrease cognitive load. (2) Empirical insights from a between-subject online study (N=117) assessing OptiCarVis against an averaged design by experts (N=8), end-users’ custom designs, and a No Visualization baseline. The study also investigated how participants interacted with OptiCarVis in three optimization conditions: a Cold-Start HITL MOBO with random initial design parameters, a Warm-Start HITL MOBO initialized by mean parameters from expert designs (N=8), and a Warm-Start HITL MOBO initialized by end-users custom parameter values. (3) Open-source [^1] Source code will be released upon acceptance implementation of a Unity-based driving simulation enabling HITL MOBO of AV functionality visualizations.
@@ 2.1. In-Vehicle Visualizations of Automated Vehicle Functionalities
Previous research evaluated different display technologies (e.g., HUDs, LED strips, and AR WSDs) for visualizing various driving-related information in AVs. For instance, Colley et al. (2020) found that the AR WSD reduced cognitive load. Besides, Häuslschmid et al. (2017) increased trust by showing the AV’s current situation interpretation of the vehicle through a miniature world or a simulated chauffeur avatar. They found the miniature world most effective, but participants’ opinions on the necessity of such visualization varied significantly. Currano et al. (2021) also tested an AR HUD and found situation awareness varied based on the driving scene’s complexity and the participants’ reported driving styles, suggesting that the HUD should be personalized and adapt to these factors. Schneider et al. (2021) evaluated explanations delivered via an AR WSD and an LED strip concerning the future vehicle trajectory. User experience increased with explanations, but adding a post-explanation via a smartphone app did not enhance it. Colley et al. (2021b) also found that an abstract visualization of objects (i.e., via a symbol with text) perceived by the AV, such as crossing dogs and children in a HUD, was sufficient.
An important factor contributing to understanding AV functionality and its safe use can be the visualization of uncertainties in detecting or predicting the driving environment and ego trajectory planning. Beller et al. (2013) focused on conveying automation uncertainty with a simple anthropomorphic symbol when system limits are reached. They observed that situational awareness and trust increased when uncertainty information was displayed. Similarly, Helldin et al. (2013) used abstract uncertainty visualization—bars indicating the system’s ability to operate—and found that users took control more quickly. However, they also found that participants trusted automation less when uncertainty was shown. As AV uncertainties on the in-vehicle dashboard may increase distractions, unnecessary glances, and cognitive load, other visualization technologies were employed. Kunze et al. (2018) used AR to visualize longitudinal and lateral control uncertainties. They found that the hue is particularly effective at conveying urgency.
However, these abstract visualizations (e.g., icons (Beller et al., 2013), ambient light (Schneider et al., 2021), or miniature worlds (Häuslschmid et al., 2017)) may not allow the user to identify the source of uncertainty. Therefore, using semantic segmentation, Colley et al. (2021a) visualized AVs’ Situation Detection. They found that their simulated AR WSD did not increase trust or cognitive load but improved situation awareness, and users rated detection-related attributes significantly better. In this context, Colley et al. (2022b) compared visualizations of the functional levels of AVs (detection, prediction, and trajectory planning) and their combinations. They found that showing the planned ego trajectory increased trust, and the prediction increased cognitive load. Flohr et al. (2023) confirmed the need to visualize AV functionalities to combat over- and undertrust in AVs by visualizing Situation Detection in an on-road wizard-of-Oz study. They found that AV functionality visualizations can increase predictability, perceived usefulness, and hedonic user experiences. Lastly, Müller et al. (2022) reused visualization designs, such as the connectivity symbol, planned trajectory, and vehicle markers. They also visualized infrastructure support via CAD and, enabled by this, displayed cars occluded behind obstacles and a gap indicator for AV merging. They found that combining all visualizations resulted in the highest trust, reliability, and understanding.
@@ 2.3. Human-in-the-Loop Multi-Objective Bayesian Optimization
HITL optimization integrates humans in its iterative parameter optimization cycles when design objectives require humans’ subjective ratings (e.g., (Chiu et al., 2020; Koyama et al., 2020; Takagi, 2001; Zhong et al., 2021)) or through performance measurements (e.g., (Dudley et al., 2019; Kadner et al., 2021; Khajah et al., 2016)).
In Human-Computer Interaction (HCI), BO has been employed in HITL optimization to tackle various design problems (Dudley et al., 2019; Kadner et al., 2021; Koyama et al., 2020; Chong et al., 2021). BO is a machine learning method for optimizing unknown and/or difficult-to-evaluate functions (Chan et al., 2022), such as black-box user models. While there are other black-box optimization methods (e.g., evolutionary and genetic algorithms, see (Alarie et al., 2021)), BO stands out due to its consistent performance (Borji and Itti, 2013) and customizability (Liao, 2023). BO iteratively evaluates and updates parameters to achieve the best results for a given objective. It balances exploration, which involves probing underexplored regions of the design space to discover potentially better designs, and exploitation, focusing on areas already identified as promising based on prior knowledge. This balance enables BO to find optimal designs with relatively few iterations, making it one of the most efficient optimization approaches (Brochu et al., 2010). Therefore, BO is well suited to the problem of AV functionality visualization design, where the relationship between design parameters and user experience is hard to model.
@@
AV functionality visualization design must consider multiple objectives like safety, trust, acceptance, predictability, and cognitive load according to the Automation Acceptance Model (Ghazizadeh et al., 2012). MOBO might be the answer to address these, as it can maximize or minimize multiple objectives simultaneously. The result is not a single optimal design but a range of solutions known as the Pareto front. This front consists of all designs in a multi-objective optimization problem that are not outperformed by any other design. Each point on the Pareto front is Pareto-optimal, meaning no other design is better in all objectives simultaneously (Marler and Arora, 2004). It illustrates the best trade-offs between conflicting objectives, where improving one (e.g., usability) may result in a decrease in another (e.g., perceived safety).
MOBO was used, for example, to support the design of touchscreen keyboards that balance speed, user familiarity, and enhanced spell-checking (Dunlop and Levine, 2012), for designing multi-finger inputs for text entry in mid-air (Sridhar et al., 2015), for creating haptic interfaces (Hayward et al., 1994), and for an interactive personalization of explanations of image classifier results (Chandramouli et al., 2023). These works demonstrated MOBO’s effectiveness in HCI design tasks, especially when studies with human participants are costly. Therefore, we argue that MOBO is an appropriate method for optimizing AV functionality visualization design. However, MOBO’s effectiveness in this domain and the associated modeling of complex end-user states, such as perceived safety, trust, and acceptance, remains unclear.
Besides, Chan et al. (2022) and Liao et al. (2023) revealed that designers felt less agency and ownership over MOBO-driven designs, even if they were of higher quality. To foster collaboration between BO-supported design approaches and designers, Koyama and Goto (2022) introduced BO as a design assistant. This allowed designers to tap into their expertise and preferences while BO offered design suggestions.
In contrast, involving end-users in AV functionality visualization design is critical, as they offer key insights into their preferences that designers cannot have. While previous research primarily explored the impact of designers in HITL MOBO design processes (and vice versa), end-user integration is left underexplored. Yet, their contributions can potentially enhance optimization efficiency and design quality compared to designers. Therefore, this work focuses on engaging end-users in the design of visualizations to address existing HITL MOBO limitations. Besides, we extend the understanding of qualities of HITL optimization established in works like (Chan et al., 2022; Liao et al., 2023) by comparing end-user-led and optimizer-driven processes.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713514isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing Empirical studies in visualizationccs: Human-centered computing Empirical studies in HCI
@@
The figure shows the OptiCarVis flow chart. The Bayesian Optimization process is applied to the results, repeatedly evaluated by study participants via a questionnaire.
@@ 1. Introduction
However, user acceptance of AV technology is not guaranteed. Studies have shown that many potential users are concerned about AV reliability (Schoettle and Sivak, 2014; Kyriakidis et al., 2015), which refers to AVs’ ability to perform driving tasks consistently safe across different situations. Both undertrust (e.g., leading to not using AVs) and overtrust (e.g., inadequately supervising the operation) present challenges for AV usage if the user’s trust is inappropriate to the actual AV reliability (Lee and See, 2004). Therefore, prior works have shown that visualizations of AV functionality are a way to enhance user experience (Colley et al., 2022b; Flohr et al., 2023; Schneider et al., 2021). Suggested visualization solutions to overcome undertrust highlighted other road users in foggy scenarios (Winter et al., 2019). In the Connected Automated Driving (CAD) context, proposals include visualizing external sensor coverage (henceforth called “CAD-covered area”) and detected road users occluded, for example, by a building (henceforth called “occluded cars”) (Müller et al., 2022). Regarding overtrust, visualizing the internal AV functionalities (e.g., Situation Detection, Situation Prediction, or Trajectory Planning) and their inherent uncertainty were evaluated (Colley et al., 2020, 2021a, 2022b; Flohr et al., 2023).
@@
However, the diversity of passengers complicates the design of AV functionality visualizations. Passengers’ subjective perceptions of safety, trust, and aesthetics differ (Mara and Meyer, 2022). In addition, passengers’ understanding of AVs’ internal functionality depends on individual knowledge and attitudes towards technology (Zhang et al., 2023). Thus, to support user understanding, designers must balance various design objectives within a complex design space (i.e., the set of possible design parameter values), for example, determining the size, transparency, and necessity of visualization elements. Traditional design methods relied on the user-centered design process (ISO 9241 (International Organization for Standardization, 2019)), standards (e.g., ISO 15005 (International Organization for Standardization, 2017)), and guidelines (e.g., the JAMA Guidelines for In-vehicle Display System (Yukinobu Nakamura, 2008)) but also experience, trial and error, and intuition especially when designing for novel AV experiences, resulting in resource-intensive user evaluations.
Prior research has explored personalization for in-vehicle displays, allowing passengers to manually adjust elements such as icon size, location, and color (Normark, 2015). However, translating personal preferences into effective design parameters can be difficult without design expertise. To address this, prior works introduced computational methods that rely on weighted user ratings to select from predefined visualization options (Yunuo et al., 2022; Zhong et al., 2023). Although these approaches measure usability, they omit perceived safety and trust, essential for fostering user acceptance (Adnan et al., 2018), and restrict the design space to a small set of discrete configurations, potentially overlooking designs that better meet passengers’ needs and preferences.
@@
To improve user experience, involving humans in the design process is important to align with their needs and preferences. A Human-in-the-Loop (HITL) process (Chan et al., 2022) can iteratively present design variants to users, using their feedback to optimize design parameters. However, this optimization is challenging due to multiple objectives, such as increasing perceived safety and trust. In non-automotive User Interface (UI) domains, (Multi-Objective) Bayesian Optimization (BO and MOBO) approaches have emerged to address these design optimization problems (Chan et al., 2022; Liao et al., 2023; Chandramouli et al., 2023; Koyama and Goto, 2022; Kadner et al., 2021). BO optimizes designs by predicting which parameter changes best meet the objectives. In MOBO–an extension of BO that can handle multiple objectives–manages trade-offs among conflicting objectives by identifying points on the Pareto front, where any improvement in one objective (e.g., perceived safety through more visualizations) would lead to another objective’s deterioration (e.g., increasing cognitive load). Such designs are termed Pareto-optimal (Marler and Arora, 2004).
@@
We demonstrate OptiCarVis by designing visualizations for an AV’s functional levels–Situation Detection (semantic segmentation), Situation Prediction (pedestrian intention icons and road user trajectories), and Trajectory Planning (AV trajectory)–along with CAD-covered areas, occluded cars, and general information (AV speed, destination, current time) to convey relevant information during automated driving stages (Colley et al., 2022b; Müller et al., 2022; Flohr et al., 2023). MOBO estimates the visualizations’ design parameter values within the design space in each HITL iteration (see Figure 1). End-users experience the selected design and provide subjective ratings after the ride. MOBO then uses these ratings to assess objective attainment and optimizes parameters for the next iteration, repeating until objectives are achieved (e.g., maximum trust ratings).
@@
To quantify optimization effectiveness, end-users evaluated their final designs based on cognitive load, predictability, trust, perceived safety, usefulness, satisfaction, and aesthetics. We compared HITL MOBO results to a baseline created by automotive UI experts (N=8) and employed webcam-based eye-tracking to monitor user reactions and attention throughout the study. In a between-subject online study with 117 participants, OptiCarVis generated personalized visualization parameters for AVs’ Situation Detection, Situation Prediction, and Trajectory Planning. These personalized designs significantly enhanced perceived safety compared to custom designs, improved predictability over expert designs, and increased trust, usefulness, and satisfaction compared to both. Participants reported satisfaction with the design process and a sense of involvement. Many deemed the design optimal but suggested incorporating additional visualization elements and expanding driving scenarios.
By leveraging OptiCarVis, automotive UI designers and end-users can navigate complex design spaces, potentially resulting in more passenger-centric UIs that could significantly increase their perceived safety, trust, and acceptance of AVs. Moreover, integrating end-users’ feedback into the design could inspire the development of more cooperative and effective end-user optimization methods that are implicitly integrated into future automotive UIs.
Contribution Statement: (1) OptiCarVis—the computational optimization of AV functionality visualization design using HITL MOBO to improve end-user trust, perceived safety, acceptance, and understanding of AVs and decrease cognitive load. (2) Empirical insights from a between-subject online study (N=117) evaluating OptiCarVis against an averaged design by experts (N=8), end-user custom designs, and a No Visualization baseline. The study also examined interactions with OptiCarVis under three optimization conditions: Cold-Start HITL MOBO with random initial parameters, Warm-Start HITL MOBO initialized with expert (N=8) mean parameters, and Warm-Start HITL MOBO initialized with end-user custom parameters. (3) Open-source implementation of a Unity-based simulation enabling HITL MOBO [^1] https://github.com/Pascal-Jansen/Bayesian-Optimization-for-Unity of AV functionality visualizations.
@@ 2.1. In-Vehicle Visualizations of Automated Vehicle Functionalities
Various display technologies (e.g., HUDs, LED strips, and AR WSDs) have been evaluated for visualizing AV information. For instance, highlighting of road users in an AR WSD can reduce cognitive load (Colley et al., 2020), while a “miniature world” view can enhance trust (Häuslschmid et al., 2017), though users’ opinion differ on its necessity. AR HUDs also affect situation awareness based on scene complexity and user driving styles, highlighting the need for personalization (Currano et al., 2021). Explaining future trajectories via an AR WSD or LED strip improves user experience, but adding a post-explanation on a smartphone confers no additional benefits (Schneider et al., 2021). Additionally, an abstract HUD view (e.g., a symbolic icon with text) can sufficiently convey critical information (e.g., crossing children) (Colley et al., 2021b).
An important factor contributing to understanding AV functionality and its safe use can be the visualization of uncertainties in detecting or predicting the driving environment and ego trajectory planning. Anthropomorphic or abstract uncertainty indicators have been shown to boost situational awareness and trust, though they may also reduce trust if the automation appears less reliable (Beller et al., 2013; Helldin et al., 2013). Due to potential increases in distraction or cognitive load, AR-based approaches have been explored instead. For instance, AR visualizations of longitudinal and lateral control uncertainties effectively convey urgency through color hues (Kunze et al., 2018).
Abstract visualizations (e.g., icons (Beller et al., 2013), ambient light (Schneider et al., 2021), miniature worlds (Häuslschmid et al., 2017)) often obscure the source of system uncertainty. In contrast, semantic segmentation of an AV’s detection process can improve situation awareness without affecting trust or cognitive load (Colley et al., 2021a). Visualizing trajectory planning can boost trust but may raise cognitive load (Colley et al., 2022b), and on-road studies confirm that showing AV functionality improves predictability, perceived usefulness, and hedonic user experiences (Flohr et al., 2023). Finally, combining established designs (e.g., planned trajectories, connectivity symbols) with infrastructure support, showing occluded vehicles and merging gaps, yields the highest trust, reliability, and understanding (Müller et al., 2022).
@@ 2.3. Human-in-the-Loop Multi-Objective Bayesian Optimization
HITL optimization integrates humans in its iterative parameter optimization cycles when design objectives require humans’ subjective ratings (e.g., (Chiu et al., 2020; Koyama et al., 2020; Takagi, 2001; Zhong et al., 2021)) or through performance measurements (e.g., (Dudley et al., 2019; Kadner et al., 2021; Khajah et al., 2016)). In Human-Computer Interaction (HCI), BO has been employed in HITL optimization to tackle various design problems (Dudley et al., 2019; Kadner et al., 2021; Koyama et al., 2020; Chong et al., 2021). BO is a machine learning method for optimizing unknown and/or difficult-to-evaluate functions (Chan et al., 2022), such as black-box user models. While there are other black-box optimization methods (e.g., evolutionary and genetic algorithms, see (Alarie et al., 2021)), BO stands out due to its consistent performance (Borji and Itti, 2013) and customizability (Liao, 2023). BO iteratively evaluates and updates parameters to achieve the best results for a given objective. It balances exploration, which involves probing underexplored regions of the design space to discover potentially better designs, and exploitation, focusing on areas already identified as promising based on prior knowledge. This balance enables BO to find optimal designs with relatively few iterations, making it one of the most efficient optimization approaches (Brochu et al., 2010). Therefore, BO is well suited to the problem of AV functionality visualization design, where the relationship between design parameters and user experience is hard to model.
@@
AV functionality visualization design must consider multiple objectives like safety, trust, acceptance, predictability, and cognitive load according to the Automation Acceptance Model (Ghazizadeh et al., 2012). MOBO might be the answer to address these, as it can maximize or minimize multiple objectives simultaneously. The result is not a single optimal design but a range of solutions known as the Pareto front. This front consists of all designs in a multi-objective optimization problem that are not outperformed by any other design. Each point on this front is Pareto-optimal, meaning no other design is better in all objectives simultaneously (Marler and Arora, 2004). It illustrates the best trade-offs between conflicting objectives, where improving one (e.g., usability) may result in a decrease in another (e.g., perceived safety).
MOBO has been used to design touchscreen keyboards balancing speed, familiarity, and spell-checking (Dunlop and Levine, 2012), multi-finger mid-air text entry (Sridhar et al., 2015), haptic interfaces (Hayward et al., 1994), and interactive personalization of image classifier explanations (Chandramouli et al., 2023). These studies demonstrate MOBO’s effectiveness in HCI design tasks, especially when human participant studies are costly. Therefore, we argue that MOBO is suitable for optimizing AV functionality visualization design. However, its effectiveness in this domain remains unclear, particularly in modeling complex end-user states such as perceived safety, trust, and acceptance.
Besides, Chan et al. (2022) and Liao et al. (2023) revealed that designers felt less agency and ownership over MOBO-driven designs, even if they were of higher quality. To foster collaboration between BO-supported design approaches and designers, Koyama and Goto (2022) introduced BO as a design assistant. This allowed designers to leverage expertise and preferences while BO offered design suggestions.
In contrast, involving end-users in AV functionality visualization design is critical, as they offer key insights into their preferences that designers cannot have. While prior research primarily explored the impact of designers in HITL MOBO design processes (and vice versa), end-user integration is left underexplored. Yet, their contributions may enhance optimization efficiency and design quality compared to designers. Therefore, this work focuses on engaging end-users in the design of visualizations to address existing HITL MOBO limitations. Besides, we extend the understanding of qualities of HITL optimization established in works like (Chan et al., 2022; Liao et al., 2023) by comparing end-user-led and optimizer-driven processes.